OpenStudy (anonymous):

What is the equation of a line in point slope form that has a slope of -2 and contains the point (1,2)

OpenStudy (compassionate):

Point slope form is y - y1 = m(x - x1) Were y1 = 2 And x1 = 1 And m = -2

OpenStudy (teddyiswatshecallsme):

y - 2 = -2(x - x) since there is an implied 1 that would just be x lol

OpenStudy (teddyiswatshecallsme):

That may be incorrect idk...

OpenStudy (compassionate):

i. Do not give out answers ii. It is always assumed that 1 = 1, but this is an equation, not an expression.

OpenStudy (teddyiswatshecallsme):

eh **** the world, the government too, tehy all just killers wearing political suits.....